OpenStudy (anonymous):

how many solutions are there to the equation 11x + 25 = 7x + 15

OpenStudy (dan815):

solve for x

OpenStudy (compassionate):

Hi, 11x + 25 = 7x + 15 Tranpose 11x - 7x = -25 + 15 4x = -10 Now divide 4 by -10.

OpenStudy (kc_kennylau):

Linear equation -> one solution

OpenStudy (kc_kennylau):

Welcome to OpenStudy :D You may want to read the code of conduct ( http://openstudy.com/code-of-conduct ). Don't forget to click on the button called "Best Answer" to give a medal to the person who helped you :)